OpenStudy (anonymous):

What does it mean to refer to a 20-in TV set or a 25-in TV set? Such units refer to the diagonal of the screen. A 25-in tv set also has the width of 20 inches. What is its height? The height of a 25-in tv is __ inches.

OpenStudy (anonymous):

20^2 + h^2 = 25^2 h^2 = 25^2 - 20^2 = 225 \[\sqrt{h} = \sqrt{225}\] h = 15